OpenStudy (anonymous):

1/sinx-1/cscx=cscx-sinx

OpenStudy (anonymous):

1/sin(x) = csc(x). csc(x) = 1/sin(x) by definition.

OpenStudy (anonymous):

Therefore 1/csc(x) = sin(x).

hero (hero):

1/sinx-1/cscx=cscx-sinx 1/sinx - sinx = 1/sinx - sinx true

OpenStudy (anonymous):

Thank you!